OpenStudy (anonymous):

please help asap! medal + fan question in the comments I don't know what I am doing wrong in this numerical proof, should I try using negative numbers instead?

Vocaloid (vocaloid):

probably an algebra mistake somewhere. it shouldn't matter what the numbers you choos

OpenStudy (anonymous):

Could you try and see if you find it because I can't seem to see what I am doing wrong, I simply followed step-by-step what my teacher did and somehow I messed up somewhere.

Vocaloid (vocaloid):

I think I see it (y+x)^2 + (ax+b) = y^2 + 2xy + x^2 + ax + b there shouldn't have been a 2 in front of the y^2

Vocaloid (vocaloid):

also, you forgot the x^2 term

OpenStudy (anonymous):

Well here is my algebraic proof maybe that will help here give me a sec.

Vocaloid (vocaloid):

try (y+x)^2 + (ax+b) = y^2 + 2xy + x^2 + ax + b and tag me if the numbers don't work

Vocaloid (vocaloid):

here's where you made your mistake:

Vocaloid (vocaloid):

(y+x)(y+x) y^2+yx+yx+x^2 ^ this part is correct but in the next step you wrote: y^x + 2yx + y^2 you changed the x^2 at the end to y^2 do you understand?

OpenStudy (anonymous):

Oh Woooooooow I hadn't even noticed that at all and I went over it 10 times. I think I need a nap, okay well thank you so much!!!!!!